In the market for tourist e-book guides, the aggregate demand and aggregate supply are given by the equations:

Qd= 2000 – 400P + 0.5I +150Pb

Qs= -516 + 300P – 60W

 

where Qd and Qs is quantity of tourist e-book guides demanded and supplied, Pis the price of a tourist e-book guide, I is household income, W is wage rate paid to tourist book guide laborers, and Pb is the price of a hardbound tourist bookguide. Assume I is Php 2,300, W is Php10, and Pb is Php 21.40. Determine the equilibrium price and quantity of tourist e-book guides in this local market. Show your solution.



1
Expert's answer
2022-03-25T15:07:28-0400

Qd=2000-400P+0.5×2,300+150×21.40=6360-400P

Qs=-516+300P-60×10=300P-1116

6360-400P=300P-1116;

7476=700P;

Equilibrium price is P=10.68

Equilibrium quantity is Q=300×10.68-1116=2088